### v2

@@ V.2 Universality Class
When it comes to universality class arguments, our results show that even a small amount of long-range interaction is sufficient to drive the phase transition into the mean-field universality class. In Fig. 7, we see that for all $p>0$ the result is consistent with a mean-field transition, and there is no significant difference between the $p=0.02$ and $p=1$ cases. While all points are consistently above the expected value of $U_{c}=0.27$, this is an effect of the finite size of the network, and Fig. 4 shows that increasing system size leads to slightly lower values of critical Binder cumulant (see the plots for $p=0.02$). This is due to a crossover between the nearest-neighbor lattice and the mean-field regimes, such that smaller lattices are not significantly impacted by rare long-range interactions. In the Binder cumulant plots, this makes the apparent value of the critical Binder cumulant closer to the 2-D Ising value for smaller systems. In terms of correlation length, the crossover is also noticeable, since we observe that for $0.02\leq p\leq 0.2$ and $L=16$ and $32$, correlation lengths were about $20-40\%$ of $L$, but only $5-15\%$ of $L$ for $L=64$ and $128$. The finite-size scaling analysis of the Binder cumulant results in a critical exponent $\nu=1$ for all values of $p$, and the expected values are $\nu=1$ for the 2-D Ising universality class ($p=0$), and $\nu=1/2$ for the mean-field universality class ($p>0$). While this may seem like a discrepancy, defining a length scale for mean-field and rewired networks is not straightforward [39]. Considering that the number of sites $N$ is the thermodynamic volume of the system, a suitable definition for the linear size of the system is $\ell=N^{1/4}$, where $4$ is the upper critical dimension [39, 40]. With that definition, our critical exponent for $p>0$ would be determined, following Eq. (10), by $\Delta U\propto\ell^{1/\nu}=L^{1/(2\nu)}$, such that our results for the appropriate exponent $\nu$ are 1/($21/(2\nu)=1\Rightarrow\nu=1/$ 2, consistent with the mean-field universality class. The finite-size scaling of the order parameter, shown in Fig. 10, strengthens the finding that $p>0$ causes the system to be in the mean-field universality class, with a clear collapse of the scaling function when an exponent $\beta=1/2$ is used, as opposed to $\beta=1/8$ for the regular lattice case. The decay of the two-point correlation function in Fig. 14 also supports this finding. Such change in universality class had been observed for similar systems in quenched small-world networks [11, 13, 14].
@@
Finally, an analysis of the correlations in time and space offers some insights on the change of universality class. The significant decrease in autocorrelation time indicates that perturbations to the stationary state dissipate quickly across the entire network, which is a direct effect of the small diameter of small-worlds. Focusing on spatial correlations, the introduction of a small fraction of long-range interactions is enough to break down system-wide correlations, as seen in Figs. 13 and 14.

### v3

@@ V.2 Universality Class
When it comes to universality class arguments, our results show that even a small amount of long-range interaction is sufficient to drive the phase transition into the mean-field universality class. In Fig. 6, we see that for all $p>0$ the result is consistent with a mean-field transition, and there is no significant difference between the $p=0.02$ and $p=1$ cases. While all points are consistently above the expected value of $U_{c}=0.27$, this is an effect of the finite size of the network, and Fig. 4 shows that increasing system size leads to slightly lower values of critical Binder cumulant (see the plots for $p=0.02$). This is due to a crossover between the nearest-neighbor lattice and the mean-field regimes, such that smaller lattices are not significantly impacted by rare long-range interactions. In the Binder cumulant plots, this makes the apparent value of the critical Binder cumulant closer to the 2-D Ising value for smaller systems. In terms of correlation length, the crossover is also noticeable, since we observe that for $0.02\leq p\leq 0.2$ and $L=16$ and $32$, correlation lengths were about $20-40\%$ of $L$, but only $5-15\%$ of $L$ for $L=64$ and $128$. The finite-size scaling analysis of the Binder cumulant results in a critical exponent $\nu=1$ for all values of $p$, and the expected values are $\nu=1$ for the 2-D Ising universality class ($p=0$), and $\nu=1/2$ for the mean-field universality class ($p>0$). While this may seem like a discrepancy, defining a length scale for mean-field and rewired networks is not straightforward [41]. Considering that the number of sites $N$ is the thermodynamic volume of the system, a suitable definition for the linear size of the system is $\ell=N^{1/4}$, where $4$ is the upper critical dimension [42, 43]. Using the upper critical dimension is known to yield the correct critical exponents for the complete graph, which is the typical network representation of the mean field case [42]. For small-world networks, the situation is similar in the sense that average separation between nodes remains finite as $N\rightarrow\infty$. With that definition, our critical exponent for $p>0$ would be determined, following Eq. (10), by $\Delta U\propto\ell^{1/\nu}=L^{1/(2\nu)}$, such that our results for the appropriate exponent $\nu$ are 1/($21/(2\nu)=1\Rightarrow\nu=1/$ 2, consistent with the mean-field universality class. The finite-size scaling of the order parameter, shown in Fig. 8, strengthens the finding that $p>0$ causes the system to be in the mean-field universality class, with a clear collapse of the scaling function when an exponent $\beta=1/2$ is used, as opposed to $\beta=1/8$ for the regular lattice case. The decay of the two-point correlation function in Fig. 12 also supports this finding. Such change in universality class had been observed for similar systems in quenched small-world networks [11, 13, 14].
@@
Finally, an analysis of the correlations in time and space offers some insights on the change of universality class. The significant decrease in autocorrelation time indicates that perturbations to the stationary state dissipate quickly across the entire network, which is a direct effect of the small diameter of small-worlds. Focusing on spatial correlations, the introduction of a small fraction of long-range interactions is enough to break down system-wide correlations, as seen in Figs. 11 and 12.